OpenStudy (anonymous):

y+3=0 need points to graph this

OpenStudy (anonymous):

y+3=0 -3 -3 y=-3 There will be a vertical line on the x-axis, on the point -3.

OpenStudy (anonymous):

so its just one point

OpenStudy (anonymous):

Its a single line. I will draw an example.

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):



OpenStudy (anonymous):

Sorry I am on a laptop.

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

You're Welcome!